6.4.15-E Assigned Media Ξ Question Help Suppose that an airline uses a seat width of 16.8 in. Assume men h ve hip breadths that are normally distributed with a mean of 14.9 in. and a standard deviation of 1 in. Cumplete parts (a) through (c) below (a) Find the probability that if an individual man is randomly selected, his hip breadth will be greater than 16.8 in. The probability is Round to four decimal places as needed.) Enter your answer in the answer box and then click Check Answer Clear All Check An * Previous Next MacBook Air F7 FB F9 4 5 7 0 0
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 14.9
std deviation   =σ= 1.0000
probability = P(X>16.8) = P(Z>1.9)= 1-P(Z<1.9)= 1-0.9713= 0.0287

( please revert if there are other parts)